CARS OBSERVATION AND ANALYSIS OF THE ν_{1} BAND OF ^{32}S^{16}O_{3}

\underline{ENGELENE t. H. CHRYSOSTOM}, TONY MASIELLO, JEFFREY BARBER, NICOLAE VUL-PANOVICI, JOSEPH W. NIBLER, Department of Chemistry, Oregon State University, 153 Gilbert Hall, Corvallis, OR 97331-4003; ALFONS WEBER, National Science Foundation, 4201 Wilson Boulevard, Arlington, VA 22230; ARTHUR MAKI, 15012 24 th Ave. S. E. Mill Creek, WA 98012; THOMAS A. BLAKE, ROBERT L. SAMS, Pacific Northwest National Laboratory, P. O. Box 999, Mail Stop K8-88, 3020 Q Avenue, Richland, WA 99352.

Sulfur trioxide is an important participant in reactions in the upper atmosphere and also in a number of industrial processes. It is a D_{3h} planar oblate top whose spectroscopy is incomplete, perhaps in part due to its corrosive properties. We are engaged in a comprehensive investigation of the fundamental and combination-overtone bands of ^{32}S^{16}O_{3} as well as of the ^{34}S^{16}O_{3}, ^{32}S^{18}O_{3} and ^{34}S^{18}O_{3} isotopic forms. High resolution (0.001 cm^{−1}) coherent anti-Stokes Raman scattering (CARS) was used at Oregon State University to determine for the first time the Q-branch structure of the IR-inactive ν_{1} symmetric stretching mode of ^{32}S^{16}O_{3} and its various isotopomers. The ν_{1} spectrum of ^{32}S^{16}O_{3} reveals two intense Q-branch regions, with surprisingly complex vibrational-rotational structure. The modeling of this has involved a subtle combination of Fermi resonance and indirect Coriolis interactions with nearby hidden states; 2ν_{4}(ℓ = 0, ±2), ν_{2} + ν_{4}(ℓ = ±1), 2ν_{2}(ℓ = 0). The analysis of the perturbed ν_{1} spectrum was made possible by locating some of these states via concurrent infrared hot-band studies at PNNL by T.A. Blake et al. The results of this combined effort will be presented for ν_{1} of ^{32}S^{16}O_{3}.